OpenStudy (abbycross167):

What is the equation for the points (3,34) and (5,39)? can someone please help me??

OpenStudy (anonymous):

ello love :D

OpenStudy (abbycross167):

howdy, can you help me?

OpenStudy (anonymous):

why yes i can... haha what type of equation do you need?

OpenStudy (abbycross167):

umm I guess one for a graph, like this equation ~Y = -1x + 3

OpenStudy (anonymous):

so slope intercept? okay this is fairly easy :) do you know where to start?

OpenStudy (abbycross167):

like this right ~ m = (y2 – y1) / (x2 – x1) ?

OpenStudy (anonymous):

yes :D we find the slope, so find the slope...

OpenStudy (abbycross167):

this is the step that I don't know how to do it

OpenStudy (anonymous):

okay so our points are (3,34) and (5,39) so we know that they are always (x,y) and we have two points, so we label them (x1,y1) and (x2,y2) so: x1 and y1 are the first point, and x2,y2 are our second point, understand?

OpenStudy (abbycross167):

ok so it would be m = (3 – 5) / (34 – 39) ??

OpenStudy (anonymous):

close 39-34/5-3 :)

OpenStudy (abbycross167):

oh ok

OpenStudy (anonymous):

yes so what is our slope?

OpenStudy (abbycross167):

5/2 ?

OpenStudy (anonymous):

yes :D good job so that is our slope, or m, so now we plug that into: y=mx+b y=5/2x+b so now we plug in one point for x and y, give it a try..

OpenStudy (abbycross167):

umm I'm not really sure but my guess is 2.5?

OpenStudy (anonymous):

uhh... well we plug in: 34=5/2(3)+b so we want to solve for b

OpenStudy (abbycross167):

b=7.5?

OpenStudy (anonymous):

no so you got to 7.5: 34=7.5+b subtract 7.5 from both sides

OpenStudy (anonymous):

34-7.5?

OpenStudy (abbycross167):

26.5?

OpenStudy (anonymous):

yes :) so our equation is y=5/2x+26.5

OpenStudy (abbycross167):

really? ok, if I show you an example from my lesson then replace the numbers in the example to the numbers in this question could you tell me if I did it right?

OpenStudy (anonymous):

sure, maybe? haha ill try

OpenStudy (abbycross167):

EXAMPLE FROM LESSON~ First find the slope (m). I am going to use the points (-2, 5) and (3,0) m = (y2 – y1) / (x2 – x1) m = (0 – 5) / (3 – -2) m = -5 /5 m = -1 Now we need to find the y intercept (b). y = mx + b 0 = (-1)(3) + b 0 = -3 + b 0 + 3 = -3 + 3 + b 3 = b For your final equation, take the value of m and the value of b and put them in the equation y = mx + b. This is your equation. Y = -1x + 3 y = -1(11) + 3 y = -11+3 y = -8 MY QUESTION / ANSWER~ m = (y2 – y1) / (x2 – x1) m = (39-34) / (5-3) m = 5/2 b=26.5 Equation = y=5/2x+26.5

OpenStudy (abbycross167):

why wasn't my answer as long as the example?

OpenStudy (anonymous):

cause we didnt really show our work, do you want me to type each step?

OpenStudy (abbycross167):

umm would you mind doing that? if you have something better to do than I'm fine with what we got, but if you showed me I think that would help me understand how to do this

OpenStudy (anonymous):

alright hold on give me a minute :)

OpenStudy (abbycross167):

ok thank you soo much I really appreciate your help!!

OpenStudy (anonymous):

we find slope: m= (39-34) / (5-3) m=5/2 now we find the y-intercept: m=5/2 using point (3,34) y=mx+b 34=5/2(3)+b 34=7.5+b 34-7.5=7.5-7.5+b 26.5=b final equation: y=mx+b y=5/2x+26.5

OpenStudy (anonymous):

there you go :)

OpenStudy (abbycross167):

ok!! you're amazing!!

OpenStudy (anonymous):

no...... you're amazing :D lol

OpenStudy (abbycross167):

haha I really appreciate your help

OpenStudy (anonymous):

its no problem, thats why im here :)